Alfred is three times as old as his brother.Six years ago, the sum of their ages was equal to alfreds present age.How old are alfred and his brother.
Answer by ankor@dixie-net.com(22740)

let a = A's age
let b = brother's age
:
Write an equation for each statement.
:
"Alfred is three times as old as his brother."
a = 3b
:
"Six years ago, the sum of their ages was equal to Alfred's present age."
(a-6)+(b-6) = a
a + b - 12 = a
a - a + b = 12
b = 12 yrs is brothers age
then
3(12) = 36 is A's age
:
:
See if that works in the 2nd sentence
6 yrs ago, their ages:
30 + 6 = 36